Understanding and Specifying Process Conditions

Good judgement comes from experience, and experience comes from bad judgement

Process Conditions

Barry LePatner A degree-of-freedom analysis of any unit operation suggests that there is flexibility in choosing some operating variables (temperature, pressure, stream composition etc.)

Stream specifications and process conditions are influenced by physical processes as well as ecomomic considerations and are not chosen arbitrarily.

The conditions used in a process most often represent an economic compromise between process performance and the capital and operating costs of the process equipment.

A degree-of-freedom analysis of any unit operation suggests that there is flexibility in choosing some operating variables (temperature, pressure, stream composition etc.)

Stream specifications and process conditions are influenced by physical processes as well as ecomomic considerations and are not chosen arbitrarily.

The conditions used in a process most often represent an economic compromise between process performance and the capital and operating costs of the process equipment.

A degree-of-freedom analysis of any unit operation suggests that there is flexibility in choosing some operating variables (temperature, pressure, stream composition etc.)

Stream specifications and process conditions are influenced by physical processes as well as ecomomic considerations and are not chosen arbitrarily.

The conditions used in a process most often represent an economic compromise between process performance and the capital and operating costs of the process equipment.

Separator and Reactor Operation

Process streams are rarely available at conditions most suitable for reactor and separator units.

Temperature, pressure and/or stream composition must be adjusted for effective process performance.

It is easier to adjust temperature and pressure of a process stream than it is to change its composition.

A decision to operate outside the pressure range of 1 to 10 bar must be justified.

A decision to operate above 400 ^{o}C must be justified.

A decision to operate outside the temperature range 40 oC to 260 ^{o}C requires special heating/cooling media, and thus must be justified.

Process streams are rarely available at conditions most suitable for reactor and separator units.

Temperature, pressure and/or stream composition must be adjusted for effective process performance.

It is easier to adjust temperature and pressure of a process stream than it is to change its composition.

A decision to operate outside the pressure range of 1 to 10 bar must be justified.

A decision to operate above 400 ^{o}C must be justified.

A decision to operate outside the temperature range 40 oC to 260 ^{o}C requires special heating/cooling media, and thus must be justified.

Process streams are rarely available at conditions most suitable for reactor and separator units.

Temperature, pressure and/or stream composition must be adjusted for effective process performance.

It is easier to adjust temperature and pressure of a process stream than it is to change its composition.

A decision to operate outside the pressure range of 1 to 10 bar must be justified.

A decision to operate above 400 ^{o}C must be justified.

A decision to operate outside the temperature range 40 oC to 260 ^{o}C requires special heating/cooling media, and thus must be justified.

Process streams are rarely available at conditions most suitable for reactor and separator units.

Temperature, pressure and/or stream composition must be adjusted for effective process performance.

It is easier to adjust temperature and pressure of a process stream than it is to change its composition.

A decision to operate outside the pressure range of 1 to 10 bar must be justified.

A decision to operate above 400 ^{o}C must be justified.

A decision to operate outside the temperature range 40 oC to 260 ^{o}C requires special heating/cooling media, and thus must be justified.

Process streams are rarely available at conditions most suitable for reactor and separator units.

Temperature, pressure and/or stream composition must be adjusted for effective process performance.

It is easier to adjust temperature and pressure of a process stream than it is to change its composition.

A decision to operate outside the pressure range of 1 to 10 bar must be justified.

A decision to operate above 400 ^{o}C must be justified.

A decision to operate outside the temperature range 40 oC to 260 ^{o}C requires special heating/cooling media, and thus must be justified.

Process streams are rarely available at conditions most suitable for reactor and separator units.

Temperature, pressure and/or stream composition must be adjusted for effective process performance.

It is easier to adjust temperature and pressure of a process stream than it is to change its composition.

A decision to operate outside the pressure range of 1 to 10 bar must be justified.

A decision to operate above 400 ^{o}C must be justified.

A decision to operate outside the temperature range 40 oC to 260 ^{o}C requires special heating/cooling media, and thus must be justified.

Reactor Operation at High Temperature

Favorable equilibrium conversion If the reaction is endothermic and approaches equilibrium, it benefits from high temperature. If the reaction is exothermic and approaches equilibrium, it benefits from low temperature.

Increase reaction rate

k = k_{0}exp (−\frac{E}{RT} )

As temperature increases, specific reaction rate k increases.

Favorable equilibrium conversion If the reaction is endothermic and approaches equilibrium, it benefits from high temperature. If the reaction is exothermic and approaches equilibrium, it benefits from low temperature.

Increase reaction rate

k = k_{0}exp (−\frac{E}{RT} )

As temperature increases, specific reaction rate k increases.

Favorable equilibrium conversion If the reaction is endothermic and approaches equilibrium, it benefits from high temperature. If the reaction is exothermic and approaches equilibrium, it benefits from low temperature.

Increase reaction rate

k = k_{0}exp (−\frac{E}{RT} )

As temperature increases, specific reaction rate k increases.

Favorable equilibrium conversion If the reaction is endothermic and approaches equilibrium, it benefits from high temperature. If the reaction is exothermic and approaches equilibrium, it benefits from low temperature.

Increase reaction rate

k = k_{0}exp (−\frac{E}{RT} )

As temperature increases, specific reaction rate k increases.

Favorable equilibrium conversion If the reaction is endothermic and approaches equilibrium, it benefits from high temperature. If the reaction is exothermic and approaches equilibrium, it benefits from low temperature.

Increase reaction rate

k = k_{0}exp (−\frac{E}{RT} )

As temperature increases, specific reaction rate k increases.

Favorable equilibrium conversion If the reaction is endothermic and approaches equilibrium, it benefits from high temperature. If the reaction is exothermic and approaches equilibrium, it benefits from low temperature.

Increase reaction rate

k = k_{0}exp (−\frac{E}{RT} )

As temperature increases, specific reaction rate k increases.

Maintain a gas phase Many catalytic reactions require gas phase for reactants and products. If boiling points are high or pressure is high, a high temperature may be required.

Improve selectivity If competing reactions occur and different reactions have different activation energies, an appropriate temperature is needed to favor the desired reaction.

Maintain a gas phase Many catalytic reactions require gas phase for reactants and products. If boiling points are high or pressure is high, a high temperature may be required.

Improve selectivity If competing reactions occur and different reactions have different activation energies, an appropriate temperature is needed to favor the desired reaction.

Maintain a gas phase Many catalytic reactions require gas phase for reactants and products. If boiling points are high or pressure is high, a high temperature may be required.

Improve selectivity If competing reactions occur and different reactions have different activation energies, an appropriate temperature is needed to favor the desired reaction.

Maintain a gas phase Many catalytic reactions require gas phase for reactants and products. If boiling points are high or pressure is high, a high temperature may be required.

Improve selectivity If competing reactions occur and different reactions have different activation energies, an appropriate temperature is needed to favor the desired reaction.

Maintain a gas phase Many catalytic reactions require gas phase for reactants and products. If boiling points are high or pressure is high, a high temperature may be required.

Improve selectivity If competing reactions occur and different reactions have different activation energies, an appropriate temperature is needed to favor the desired reaction.

Separator Operation at Low Pressure

Obtain gas phase for vapor-liquid equilibrium This situation arises quite frequently when high boiling point materials are needed to be distilled. An example is the distillation of crude oil in which the bottom of the atmospheric column is typically operated in the region of 310 ^{o}C to 340 ^{o}C.

Temperature sensitive material This situation arises frequently in the manufacture of specialty chemicals. Examples of material that are temperature sensitive are fragrance chemicals, flavor chemicals, pharmaceuticals.

Obtain gas phase for vapor-liquid equilibrium This situation arises quite frequently when high boiling point materials are needed to be distilled. An example is the distillation of crude oil in which the bottom of the atmospheric column is typically operated in the region of 310 ^{o}C to 340 ^{o}C.

Temperature sensitive material This situation arises frequently in the manufacture of specialty chemicals. Examples of material that are temperature sensitive are fragrance chemicals, flavor chemicals, pharmaceuticals.

Obtain gas phase for vapor-liquid equilibrium This situation arises quite frequently when high boiling point materials are needed to be distilled. An example is the distillation of crude oil in which the bottom of the atmospheric column is typically operated in the region of 310 ^{o}C to 340 ^{o}C.

Temperature sensitive material This situation arises frequently in the manufacture of specialty chemicals. Examples of material that are temperature sensitive are fragrance chemicals, flavor chemicals, pharmaceuticals.

Obtain gas phase for vapor-liquid equilibrium This situation arises quite frequently when high boiling point materials are needed to be distilled. An example is the distillation of crude oil in which the bottom of the atmospheric column is typically operated in the region of 310 ^{o}C to 340 ^{o}C.

Temperature sensitive material This situation arises frequently in the manufacture of specialty chemicals. Examples of material that are temperature sensitive are fragrance chemicals, flavor chemicals, pharmaceuticals.

Obtain gas phase for vapor-liquid equilibrium This situation arises quite frequently when high boiling point materials are needed to be distilled. An example is the distillation of crude oil in which the bottom of the atmospheric column is typically operated in the region of 310 ^{o}C to 340 ^{o}C.

Temperature sensitive material This situation arises frequently in the manufacture of specialty chemicals. Examples of material that are temperature sensitive are fragrance chemicals, flavor chemicals, pharmaceuticals.

Obtain gas phase for vapor-liquid equilibrium This situation arises quite frequently when high boiling point materials are needed to be distilled. An example is the distillation of crude oil in which the bottom of the atmospheric column is typically operated in the region of 310 ^{o}C to 340 ^{o}C.

Temperature sensitive material This situation arises frequently in the manufacture of specialty chemicals. Examples of material that are temperature sensitive are fragrance chemicals, flavor chemicals, pharmaceuticals.

Product Present in Reactor Feed

Having products in the feed to the reactor usually imply a potential loss of profit. However, products may be present in the feed under the following conditions:

Product cannot be easily separated from recycled feed material.

Recycled product retards the formation of unwanted by-products formed from side reactions.

Product acts as a diluent to control the rate of reaction.

Having products in the feed to the reactor usually imply a potential loss of profit. However, products may be present in the feed under the following conditions:

Product cannot be easily separated from recycled feed material.

Recycled product retards the formation of unwanted by-products formed from side reactions.

Product acts as a diluent to control the rate of reaction.

Having products in the feed to the reactor usually imply a potential loss of profit. However, products may be present in the feed under the following conditions:

Product cannot be easily separated from recycled feed material.

Recycled product retards the formation of unwanted by-products formed from side reactions.

Product acts as a diluent to control the rate of reaction.

Reading Assignment

Read Chapter 4 from TBWS. In particular, please study the following tables in detail.

Table 4.1: Possible reasons for operating reactors and separators outside the temperature range of special concern (page 120)

Table 4.2: Possible reasons for operating reactors and separators outside the pressure range of special concern (page 121)

Table 4.3: Possible reasons for nonstoichiometric reactor feed compositions of special concern (page 122)

Table 4.4: Changes in process conditions that are of special concern for a stream passing through a single piece of equipment (page 124)

Understanding and Specifying Process Conditions – p.9/15

Read Chapter 4 from TBWS. In particular, please study the following tables in detail.

Table 4.1: Possible reasons for operating reactors and separators outside the temperature range of special concern (page 120)

Table 4.2: Possible reasons for operating reactors and separators outside the pressure range of special concern (page 121)

Table 4.3: Possible reasons for nonstoichiometric reactor feed compositions of special concern (page 122)

Table 4.4: Changes in process conditions that are of special concern for a stream passing through a single piece of equipment (page 124)

Understanding and Specifying Process Conditions – p.9/15

Read Chapter 4 from TBWS. In particular, please study the following tables in detail.

Table 4.1: Possible reasons for operating reactors and separators outside the temperature range of special concern (page 120)

Table 4.2: Possible reasons for operating reactors and separators outside the pressure range of special concern (page 121)

Table 4.3: Possible reasons for nonstoichiometric reactor feed compositions of special concern (page 122)

Table 4.4: Changes in process conditions that are of special concern for a stream passing through a single piece of equipment (page 124)

Understanding and Specifying Process Conditions – p.9/15

Read Chapter 4 from TBWS. In particular, please study the following tables in detail.

Table 4.1: Possible reasons for operating reactors and separators outside the temperature range of special concern (page 120)

Table 4.2: Possible reasons for operating reactors and separators outside the pressure range of special concern (page 121)

Table 4.3: Possible reasons for nonstoichiometric reactor feed compositions of special concern (page 122)

Table 4.4: Changes in process conditions that are of special concern for a stream passing through a single piece of equipment (page 124)

Understanding and Specifying Process Conditions – p.9/15

Heuristics in Design

It is important to be able to apply knowledge gained through experience to future problems.

Heuristics are useful in rapidly determining feasibility of a process design. A heuristic does not guarantee a solution. It may contradict other heuristics. It can reduce the time to solve a problem. It’s acceptance depends on the immediate context instead of on an absolute standard.

It is important to be able to apply knowledge gained through experience to future problems.

Heuristics are useful in rapidly determining feasibility of a process design. A heuristic does not guarantee a solution. It may contradict other heuristics. It can reduce the time to solve a problem. It’s acceptance depends on the immediate context instead of on an absolute standard.

It is important to be able to apply knowledge gained through experience to future problems.

Heuristics are useful in rapidly determining feasibility of a process design. A heuristic does not guarantee a solution. It may contradict other heuristics. It can reduce the time to solve a problem. It’s acceptance depends on the immediate context instead of on an absolute standard.

It is important to be able to apply knowledge gained through experience to future problems.

Heuristics are useful in rapidly determining feasibility of a process design. A heuristic does not guarantee a solution. It may contradict other heuristics. It can reduce the time to solve a problem. It’s acceptance depends on the immediate context instead of on an absolute standard.

Physical Property Heuristics

Units  Organic Liq. Organic Vap.
Heat Capacity kJ/kg^{o}C 1.0 − 2.5  2.0 − 4.0
Thermal Cond. W/m^{o}C 0.10 − 0.20 0.02 − 0.06
Viscosity  kg/ms wide range 10 − 30 × 10^{−6}
Prandtl No.  10 − 1000  0.7 − 0.8

Units  Organic Liq. Organic Vap.
Heat Capacity kJ/kg^{o}C 1.0 − 2.5  2.0 − 4.0
Thermal Cond. W/m^{o}C 0.10 − 0.20 0.02 − 0.06
Viscosity  kg/ms wide range 10 − 30 × 10^{−6}
Prandtl No.  10 − 1000  0.7 − 0.8

Units  Organic Liq. Organic Vap.
Heat Capacity kJ/kg^{o}C 1.0 − 2.5  2.0 − 4.0
Thermal Cond. W/m^{o}C 0.10 − 0.20 0.02 − 0.06
Viscosity  kg/ms wide range 10 − 30 × 10^{−6}
Prandtl No.  10 − 1000  0.7 − 0.8

Units  Organic Liq. Organic Vap.
Heat Capacity kJ/kg^{o}C 1.0 − 2.5  2.0 − 4.0
Thermal Cond. W/m^{o}C 0.10 − 0.20 0.02 − 0.06
Viscosity  kg/ms wide range 10 − 30 × 10^{−6}
Prandtl No.  10 − 1000  0.7 − 0.8

Units  Organic Liq. Organic Vap.
Heat Capacity kJ/kg^{o}C 1.0 − 2.5  2.0 − 4.0
Thermal Cond. W/m^{o}C 0.10 − 0.20 0.02 − 0.06
Viscosity  kg/ms wide range 10 − 30 × 10^{−6}
Prandtl No.  10 − 1000  0.7 − 0.8

Units  Organic Liq. Organic Vap.
Heat Capacity kJ/kg^{o}C 1.0 − 2.5  2.0 − 4.0
Thermal Cond. W/m^{o}C 0.10 − 0.20 0.02 − 0.06
Viscosity  kg/ms wide range 10 − 30 × 10^{−6}
Prandtl No.  10 − 1000  0.7 − 0.8

Heuristics for Distillation Towers

Distillation is usually the most economical method for separating liquids.

For multicomponent mixtures:

Perform the easiest separation first. When neither relative volatility nor feed composition vary widely, remove components one by one as overhead products. When the adjacent ordered components in the feed vary widely in relative volatility, sequence the splits in order of decreasing volatility. When the concentrations in the feed vary widely but the relative volatilities do not, remove the components in order of decreasing concentration.

Distillation is usually the most economical method for separating liquids.

For multicomponent mixtures:

Perform the easiest separation first. When neither relative volatility nor feed composition vary widely, remove components one by one as overhead products. When the adjacent ordered components in the feed vary widely in relative volatility, sequence the splits in order of decreasing volatility. When the concentrations in the feed vary widely but the relative volatilities do not, remove the components in order of decreasing concentration.

Distillation is usually the most economical method for separating liquids.

For multicomponent mixtures:

Perform the easiest separation first. When neither relative volatility nor feed composition vary widely, remove components one by one as overhead products. When the adjacent ordered components in the feed vary widely in relative volatility, sequence the splits in order of decreasing volatility. When the concentrations in the feed vary widely but the relative volatilities do not, remove the components in order of decreasing concentration.

Distillation is usually the most economical method for separating liquids.

For multicomponent mixtures:

Perform the easiest separation first. When neither relative volatility nor feed composition vary widely, remove components one by one as overhead products. When the adjacent ordered components in the feed vary widely in relative volatility, sequence the splits in order of decreasing volatility. When the concentrations in the feed vary widely but the relative volatilities do not, remove the components in order of decreasing concentration.

Distillation is usually the most economical method for separating liquids.

For multicomponent mixtures:

Perform the easiest separation first. When neither relative volatility nor feed composition vary widely, remove components one by one as overhead products. When the adjacent ordered components in the feed vary widely in relative volatility, sequence the splits in order of decreasing volatility. When the concentrations in the feed vary widely but the relative volatilities do not, remove the components in order of decreasing concentration.

Distillation is usually the most economical method for separating liquids.

For multicomponent mixtures:

Perform the easiest separation first. When neither relative volatility nor feed composition vary widely, remove components one by one as overhead products. When the adjacent ordered components in the feed vary widely in relative volatility, sequence the splits in order of decreasing volatility. When the concentrations in the feed vary widely but the relative volatilities do not, remove the components in order of decreasing concentration.

Tower operating pressure is usually determined by temperature of the condensing media or by the maximum allowable reboiler temperature.

Economical optimum reflux ratio is in the range of 1.2-1.5 times the minimum reflux ratio.

The economically optimum number of theoretical stages is near twice the minimum value.

Reflux drums are usually horizontal, with a liquid holdup of 5 minutes half full.

Tower height is limited to about 53 m.

Tower operating pressure is usually determined by temperature of the condensing media or by the maximum allowable reboiler temperature.

Economical optimum reflux ratio is in the range of 1.2-1.5 times the minimum reflux ratio.

The economically optimum number of theoretical stages is near twice the minimum value.

Reflux drums are usually horizontal, with a liquid holdup of 5 minutes half full.

Tower height is limited to about 53 m.

Tower operating pressure is usually determined by temperature of the condensing media or by the maximum allowable reboiler temperature.

Economical optimum reflux ratio is in the range of 1.2-1.5 times the minimum reflux ratio.

The economically optimum number of theoretical stages is near twice the minimum value.

Reflux drums are usually horizontal, with a liquid holdup of 5 minutes half full.

Tower height is limited to about 53 m.

Tower operating pressure is usually determined by temperature of the condensing media or by the maximum allowable reboiler temperature.

Economical optimum reflux ratio is in the range of 1.2-1.5 times the minimum reflux ratio.

The economically optimum number of theoretical stages is near twice the minimum value.

Reflux drums are usually horizontal, with a liquid holdup of 5 minutes half full.

Tower height is limited to about 53 m.

Tower operating pressure is usually determined by temperature of the condensing media or by the maximum allowable reboiler temperature.

Economical optimum reflux ratio is in the range of 1.2-1.5 times the minimum reflux ratio.

The economically optimum number of theoretical stages is near twice the minimum value.

Reflux drums are usually horizontal, with a liquid holdup of 5 minutes half full.

Tower height is limited to about 53 m.

Heuristics for Heat Exchangers

In a shell and tube heat exchanger, the tube side is for corrosive, fouling, scaling and high pressure fluids.

In a shell and tube heat exchanger, the shell side is for viscous and condensing fluids.

Pressure drops are 0.1 bar for boiling, and 0.20-0.62 bar for other services.

Minimum temperature approach is 10^{o}C for fluids and 5^{o}C for refrigerants.

Cooling water inlet is 30^{o}C and maximum outlet is 45^{o}C.

Double-pipe heat-exchanger is competitive at duties requiring 9.3-18.6 m^{2}.

Heat transfer coefficients (in W/m^{2}K) for estimating purposes:

Water to liquid 850
Condensers  850
Liquid to Liquid 280
Liquid to Gas 60
Gas to Gas  30
Reboiler  1140

Understanding and Specifying Process Conditions – p.14/15

In a shell and tube heat exchanger, the tube side is for corrosive, fouling, scaling and high pressure fluids.

In a shell and tube heat exchanger, the shell side is for viscous and condensing fluids.

Pressure drops are 0.1 bar for boiling, and 0.20-0.62 bar for other services.

Minimum temperature approach is 10^{o}C for fluids and 5^{o}C for refrigerants.

Cooling water inlet is 30^{o}C and maximum outlet is 45^{o}C.

Double-pipe heat-exchanger is competitive at duties requiring 9.3-18.6 m^{2}.

Heat transfer coefficients (in W/m^{2}K) for estimating purposes:

Water to liquid 850
Condensers  850
Liquid to Liquid 280
Liquid to Gas 60
Gas to Gas  30
Reboiler  1140

Understanding and Specifying Process Conditions – p.14/15

In a shell and tube heat exchanger, the tube side is for corrosive, fouling, scaling and high pressure fluids.

In a shell and tube heat exchanger, the shell side is for viscous and condensing fluids.

Pressure drops are 0.1 bar for boiling, and 0.20-0.62 bar for other services.

Minimum temperature approach is 10^{o}C for fluids and 5^{o}C for refrigerants.

Cooling water inlet is 30^{o}C and maximum outlet is 45^{o}C.

Double-pipe heat-exchanger is competitive at duties requiring 9.3-18.6 m^{2}.

Heat transfer coefficients (in W/m^{2}K) for estimating purposes:

Water to liquid 850
Condensers  850
Liquid to Liquid 280
Liquid to Gas 60
Gas to Gas  30
Reboiler  1140

Understanding and Specifying Process Conditions – p.14/15

In a shell and tube heat exchanger, the tube side is for corrosive, fouling, scaling and high pressure fluids.

In a shell and tube heat exchanger, the shell side is for viscous and condensing fluids.

Pressure drops are 0.1 bar for boiling, and 0.20-0.62 bar for other services.

Minimum temperature approach is 10^{o}C for fluids and 5^{o}C for refrigerants.

Cooling water inlet is 30^{o}C and maximum outlet is 45^{o}C.

Double-pipe heat-exchanger is competitive at duties requiring 9.3-18.6 m^{2}.

Heat transfer coefficients (in W/m^{2}K) for estimating purposes:

Water to liquid 850
Condensers  850
Liquid to Liquid 280
Liquid to Gas 60
Gas to Gas  30
Reboiler  1140

Understanding and Specifying Process Conditions – p.14/15

Read Chapter 9 from TBWS. In particular, please study the tables 9.2-9.18 (pages 337-352) in detail.

Read Chapter 9 from TBWS. In particular, please study the tables 9.2-9.18 (pages 337-352) in detail.